# CONCENTRATED SUPERELLIPTICAL MARKET MAKER

### v1

@@ 2.1 AMM Invariant
Uniswap v2 introduced a continuous uniform liquidity distribution [7] allowing for the trading of a token in the price range of zero to infinity. Yet prices are not infinite nor bound by the zero barrier, to address part of this problem Uniswap v3 was introduced to allow for the concentration of liquidity based on an LP’s preferred viewpoint on the range of price movement between zero and infinity [8]. The constrained nature of a discrete uniform liquidity distribution of a Uniswap v3 position creates a problem of capital inefficiency in the tails where an LP can wind up out of range while earning no fees [Figure2]. To deal with being out of range, multiple Uniswap v3 positions can be customized along curves or, alternatively, AMM invariants can be made that match particular price behavior such as Geometric Brownian Motion with Replicating Market Makers (RMM) [9].This approach is extended by making the empirical observation that real world financial assets, as they continue to migrate onto the blockchain, can exhibit power law behavior [Figure1B][10][11] as well as enter the negative price domain [12].
@@
where $a$ and $b$ are offsets equal to $k$ to pin liquidity. The liquidity distribution of this Concentrated Circular Market Maker (CCMM) is derived in Appendix B and can be extended along a Lamé curve [13] to create a Concentrated Super-Elliptical Market Maker (CSEMM) to further widen/narrow and skew the liquidity allocation with the invariant becoming:
@@
where $u(x)=\frac{ln(2)}{ln(\frac{x}{x-1})}$. Parameters $\alpha$ and $\beta$ allow for the extension of the liquidity in the left and right tail. By setting $\alpha=\beta$ and decreasing their values towards $2$ one can approach the liquidity concentration of $x+y=k$, by widening $\alpha$ and $\beta$ towards infinity one increases the tails of the distribution. At $\alpha=\beta=\frac{1}{\sqrt{\frac{\left(\sqrt{2}-1\right)^{2}}{2}}}$ we retrieve the CCMM from equation 1.
The swap functions are given in Appendix C, compared to other invariants in Figure 3, and the RMM in desmos: https://www.desmos.com/calculator/twsvm9djv5. Its unique feature allows one to trade two assets where either price can go negative between one another. While appearing bizarre, it is normal to be able to exchange a barrel of oil costing negative four dollars for two natural gas costing a negative two dollars each. This is at the present moment not doable in centralized exchanges, which in the middle of COVID-19 had difficulty accounting for negative oil prices by having to switch to Arithmetic Brownian Motion from Geometric Brownian Motion via the Bachelier Model [14]. Furthermore, for an AMM where both $x$ and $y$ can go negative, the value of the LP position [Figure 4] also loops on itself and, in the event of borrowing such an LP position, can exhibit double convexity [15].
@@ 2.2 Applications
Additionally, the purpose does not necessarily have to apply to the underlying price going negative. It could apply to an offset representation of the price. For example, an asset can be worth \$100 and yet its offset price token can be \$0 and now a negative offset price of \$ -1 would mean a reduction in the price of the underlying to \$99, creating a variety of intriguing non-linear payoffs resembling options shown in Figure 4 and desmos: https://www.desmos.com/calculator/qkexksh7vf.
@@ 3.1 A - Negative Liquidity
3.2 B - Liquidity Distribution Derivation
Start with the full invariant
@@ 3.2 B - Liquidity Distribution Derivation
$y_{x}=-\frac{dy_{n}}{dx}=\pm\frac{\ -x+a}{\sqrt{k^{2}-x^{2}+2ax-a^{2}}}.$ (9)
@@
Rewrite it as a function of price ($\sqrt{p}$) by raising $x^{2}$, rewrite it as $l(x)$, take its derivative to get the liquidity distribution in price space
@@
The distribution happens to have the tail of a Student’s-T distribution with degree of freedom $df=2$. It is also a heavy tailed distribution with a Pareto tail index of $\alpha=3$
@@
Converting it to tick space $t$ we get a liquidity fingerprint with heavier tails than the RMM’s Gaussian distribution:
$L(t)=\pm\frac{2ke^{3t}}{\left(1+e^{4t}\right)^{\frac{3}{2}}}.$ (14)
The value function [19] of an LP payoff in the positive domain is given by
$V_{c}(p)=2p\frac{a+bp-\sqrt{k^{2}p^{2}-a^{2}p^{2}+2abp+k^{2}-b^{2}}}{1+p^{2\ }}$ (15)
with the Greeks for Delta=$\Delta_{c}$, Gamma=$\Gamma_{c}$, and Theta=$\Theta_{c}$ being available in desmos at https://www.desmos.com/calculator/twsvm9djv5. Expected theta is derived from the relationship of concavity to yield as originally outlined by Louis Bachelier in 1900 [20]
@@ 3.4 D - Asymmetric Negative Liquidity
$y_{x}=-\frac{dy_{n}}{dx}=\frac{1-\sqrt{x}}{\sqrt{x}}.$ (23)
@@
First we solve it as a function of positive price ($\sqrt{p}$) by raising $x^{2}$, rewrite it as $l_{+}(x)$
@@
$L_{+}(x)=\frac{dl(x)}{dx}=\frac{4kx^{3}}{\left(1+x^{2}\right)^{3}}\{x>0\}.$ (28)
Converting it to tick space $t$ we get:
$L_{+}(t)=\frac{4ke^{3t}}{\left(1+e^{2t}\right)^{3}}\{t>0\}.$ (29)
@@
$L_{-}(x)=\frac{dl(x)}{dx}=-\frac{-4kx^{3}}{\left(1-x^{2}\right)^{3}}\{x<0\}.$ (32)
@@
$L_{-}(t)=-\frac{-4ke^{3t}}{\left(1-e^{2t}\right)^{3}}\{t<0\}.$ (33)
@@ 4 Disclaimer
Figure 2: A CPMM follows a uniform distribution (blue). The liquidity distribution of an RMM [18] follows a Gaussian distribution (black), yet financial assets have the tendency to exhibit fat tail behavior which can cause an LP to under-allocate liquidity in the tails (red area in 2B).
@@
Figure 4: LP Payoffs of CCMM. Note that an LP Position with negative prices can exhibit convexity. This is intuitive. Consider an LP who is willing to receive payment in token 1 (USDC) for receiving token 2 (OIL). For example, an oil trader may receive payment for taking oil inventory during COVID-19 in 2020, yet as soon as the market rebounds, the oil trader proceeds to sell out of appreciated oil back into USDC automatically.
Figure 5: Unit circle with price stretching along it from negative to positive infinity. 5A: Uniswap v2 uniform liquidity in blue stretching across the positive domain and the negative domain in red with two Uniswap v3 concentrated liquidity positions in purple on the positive price domain. 5B: RMM with no negative liquidity given that price cannot go beyond zero. 5C: liquidity distribution of a CCMM with negative liquidity. 5D: liquidity distribution of a parabolic invariant with negative liquidity increasing as price enters the negative domain.

### v2

@@ 2.1 AMM Invariant
Uniswap v2 introduced a continuous uniform liquidity distribution [7] allowing for the trading of a token in the price range of zero to infinity. Yet prices are not infinite nor bound by the zero barrier, to address part of this problem Uniswap v3 was introduced to allow for the concentration of liquidity based on an LP’s preferred viewpoint on the range of price movement between zero and infinity [8]. The constrained nature of a discrete uniform liquidity distribution of a Uniswap v3 position creates a problem of capital inefficiency in the tails where an LP can wind up out of range while earning no fees [Figure2]. To deal with being out of range, multiple Uniswap v3 positions can be customized along curves or, alternatively, AMM invariants can be made that match particular price behavior such as Geometric Brownian Motion in Black Scholes with Replicating Market Makers (RMM) [9].This approach is extended by making the empirical observation that real world financial assets, as they continue to migrate onto the blockchain, can exhibit heavy tails [Figure1B][10][11] as well as enter the negative price domain [12].
@@
where $a$ and $b$ are offsets equal to $k$ to pin liquidity. The liquidity fingerprint of this Concentrated Circular Market Maker (CCMM) is derived in Appendix B and can be extended along a Lamé curve [13] to create a Concentrated Super-Elliptical Market Maker (CSEMM) to further widen/narrow and skew the liquidity allocation with the invariant becoming:
@@
where $u(x)=\frac{ln(2)}{ln(\frac{x}{x-1})}$. Parameters $\alpha$ and $\beta$ allow for the extension of the liquidity in the left and right tail. By setting $\alpha=\beta$ and decreasing their values towards $2$ one can approach the liquidity concentration of $x+y=k$, by widening $\alpha$ and $\beta$ towards infinity one increases the tails of the distribution until it converges to the Logarithmic Market Scoring Rule (LMSR) invariant. At $\alpha=\beta=\frac{1}{\sqrt{\frac{\left(\sqrt{2}-1\right)^{2}}{2}}}$ we retrieve the CCMM from equation 1.
The swap functions are given in Appendix C, compared to other invariants in Figure 3 and the RMM with the LMSR in desmos: https://www.desmos.com/calculator/mbohmvmytg. Its unique feature allows one to trade two assets where either price can go negative between one another. While appearing bizarre, it is normal to be able to exchange a barrel of oil costing negative four dollars for two natural gas costing a negative two dollars each. This is at the present moment not doable in centralized exchanges, which in the middle of COVID-19 had difficulty accounting for negative oil prices by having to switch to Arithmetic Brownian Motion from Geometric Brownian Motion via the Bachelier Model [14]. Furthermore, for an AMM where both $x$ and $y$ can go negative, the value of the LP position also loops on itself and, in the event of borrowing such an LP position during a negative risk free rate, can exhibit double convexity [15].
@@ 2.2 Applications
Additionally, the purpose does not necessarily have to apply to the underlying price going negative. It could apply to an offset representation of the price. For example, an asset can be worth \$100 and yet its offset price token can be \$0 and now a negative offset price of \$ -1 would mean a reduction in the price of the underlying to \$99, creating a variety of intriguing non-linear payoffs [Figure4].
@@ 3.1 A - Negative Liquidity
3.2 B - Liquidity Fingerprint Derivation
By taking the second derivative with respect to the square root of price we can derive the liquidity fingerprint in price space. Start with the full invariant
@@ 3.2 B - Liquidity Fingerprint Derivation
$y_{x}=-\frac{dy}{dx}=\pm\frac{\ -x+a}{\sqrt{k^{2}-x^{2}+2ax-a^{2}}}.$ (9)
@@
Rewrite it as a function of price ($\sqrt{p}$) by raising $x^{2}$ and express it as $l(x)$
@@
Taking its derivative to get the liquidity distribution in price space. The distribution happens to have a tail with a Pareto tail index of $\alpha=3$
@@
Converting it to tick space $t$ where $t=e^{\frac{x}{2}}$ to get a liquidity fingerprint with heavier tails than the covered call Gaussian distribution:
$L(t)=\pm\frac{2ke^{\frac{3t}{2}}}{\left(1+e^{2t}\right)^{\frac{3}{2}}}.$ (14)
The value function [19] of an LP payoff is given by the Legendre transform with the Greeks for Delta=$\Delta_{c}$, Gamma=$\Gamma_{c}$, and Theta=$\Theta_{c}$ being available in desmos at https://www.desmos.com/calculator/mbohmvmytg. Expected theta is derived from the relationship of concavity to yield as originally outlined by Louis Bachelier in 1900 [20]
@@ 3.4 D - Asymmetric Negative Liquidity
$y_{x}=-\frac{dy}{dx}=\frac{1-\sqrt{x}}{\sqrt{x}}.$ (22)
@@
First we solve it as a function of positive price ($\sqrt{p}$) by raising $x^{2}$, expressing it as $l_{+}(x)$
@@
$L_{+}(x)=\frac{dl(x)}{dx}=\frac{4x^{3}}{\left(1+x^{2}\right)^{3}}\{x>0\}.$ (27)
Converting it to tick space $t$ where $t=e^{\frac{x}{2}}$ we get:
$L_{+}(t)=\frac{4e^{\frac{3t}{2}}}{\left(1+e^{t}\right)^{3}}\{t>0\}.$ (28)
@@
$L_{-}(x)=\frac{dl(x)}{dx}=-\frac{-4x^{3}}{\left(1-x^{2}\right)^{3}}\{x<0\}.$ (31)
@@
$L_{-}(t)=-\frac{-4e^{\frac{3t}{2}}}{\left(e^{t}-1\right)^{3}}\{t<0\}.$ (32)
@@ 4 Disclaimer
Figure 2: A CPMM follows a uniform distribution (blue). The liquidity distribution of a covered call RMM [18] follows a Gaussian distribution (black), yet financial assets have the tendency to exhibit fat tail behavior which can cause an LP to under-allocate liquidity in the tails (red area in 2B).
@@
Figure 4: The LP Payoff of a CCMM exhibits concavity when an invariant is constructed. It will exhibit convexity when the LP position is borrowed. The non-linear payoff can further be sharpened and skewed with a CSEMM.
Figure 5: Unit circle with price stretching along it from negative to positive infinity. 5A: Uniswap v2 uniform liquidity in blue stretching across the positive domain and the negative domain in red with two Uniswap v3 concentrated liquidity positions in purple on the positive price domain. 5B: RMM covered call with no negative liquidity given that price cannot go beyond zero. 5C: liquidity distribution of a CCMM with negative liquidity. 5D: liquidity distribution of a parabolic invariant with negative liquidity increasing as price enters the negative domain.